OpenStudy (cucacula):

evaluate ∫x ln5 xdx using integration by parts. Use U=ln 5x . anyonce can help me?

sam (.sam.):

\[\Huge \int\limits xln(5x)dx?\]

OpenStudy (cucacula):

yes @.Sam.

OpenStudy (aravindg):

integration by parts?

OpenStudy (cucacula):

Use U=ln 5x. @AravindG

OpenStudy (aravindg):

Oh sorry ..I didnt read that

sam (.sam.):

First we'll do u-sub \[ u=5x ~~~~~~~\frac{u}{5}=x\\ \\ du=5dx \\ \\ \frac{du}{5}=dx \\ \\ \frac{1}{25} \int\limits uln(u)du\]

sam (.sam.):

IBP\[Let \\ \\ J=\ln(u)~~~~~dk=udu \\ \\ dJ=\frac{1}{u}du ~~~~~~k=\frac{u^2}{2} \\ \\ \frac{1}{25}[\frac{u^2}{2}\ln(u)-\int\limits \frac{u^{\cancel 2}}{2}\frac{1}{\cancel u}du ]\\ \\ \frac{1}{25}[\frac{u^2}{2}\ln(u)-\frac{u^2}{4}] \\ \\ \frac{1}{25}[\frac{25x^2}{2}\ln(5x)-\frac{25x^2}{4}] \\ \\ \frac{x^2}{2}\ln(5x)-\frac{x^2}{4}\]

sam (.sam.):

Oh and don't forget the constant

sam (.sam.):

The formula for J and K integration by parts here is \[\int\limits xydx \\ \\ \int\limits \overbrace{x}^{J} \overbrace{ydy}^{dK} \\ \text{Find dJ and K then substitute into} \\ JK-\int\limits K~dJ\]

OpenStudy (cucacula):

how to find the dJ and k ? @.Sam.

sam (.sam.):

For your question, we set \[J=\ln(u) ~~~ \text{and} ~~~ dK=udu \\ \\ \] So differentiating J, we get dJ, and integrating dK, we get K \[ dJ=\frac{1}{u}du ~~~~~~k=\frac{u^2}{2}\]

sam (.sam.):

Then substitute into this\[JK-\int\limits\limits K~dJ\] and get \[ \frac{1}{25}[\frac{u^2}{2}\ln(u)-\int\limits\limits \frac{u^{ 2}}{2}\frac{1}{ u}du ]\]

OpenStudy (anonymous):

there is a formula to solve such types of problems that contains two terms one you can integrate easily and second that you can't. so in order to solve such problems the formula is (first term) cross (integration of second term)-integration{(differentiation of first term)-(integration of second term)dx}dx firs term=that you can't integrate easily second term= that you can integrate easily your first term is ln(5x) and second term is x.